OpenStudy (anonymous):

a bag contains 3white and 4red balls.two draws of one ball each are made without replacement.what is the probability that one is red and other is white

OpenStudy (raden):

Probability one is red of the first draw is 4/7 Probability one is white of the 2nd draw is 3/6 =1/2 multiply of both up, we get 4/7 * 1/2 = 4/14 = 2/7

OpenStudy (raden):

does that make sense, @pooja123 ?

OpenStudy (anonymous):

the answer is 4/7.

OpenStudy (raden):

4/7 is just for the probalility of one red only :) do u want draw 2 balls, right ?

OpenStudy (anonymous):

yes they used the formula like 2into4into3/7into 6

OpenStudy (raden):

hmm. i dont understand that, actually the formula to get the probability of an event using : P(A) = number of event A / number all event

OpenStudy (nubeer):

hmm well she is right the answer would be 4/7

OpenStudy (raden):

why, can u explain more, nubber ?

OpenStudy (nubeer):

we found the probability 2/7 when first red and second white so P[RW] = 2/7 this is one case another possibility is that first is white and second is red P[WR] = 3/7 * 4/6 = 2/7 so total probability of this event = p[WR]+p[RW]

OpenStudy (anonymous):

thanks

OpenStudy (raden):

wait, is draw it one by one Or 2 balls direct ?

OpenStudy (nubeer):

hmm i think one by one...

OpenStudy (raden):

hmmmm.. yeah, i see now. the question doesnt says the the first draw must red or white, thanks for your corrections @nubeer and sorry for @pooja123 :)

OpenStudy (nubeer):

glad to help :)

OpenStudy (anonymous):

thank u guys

OpenStudy (raden):

im not becareful with this question :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can u see my isosceles triangle solution for one sec